# Benchmarking Language Model Creativity: A Case Study on Code Generation

### v1

@@ Abstract
As LLMs become increasingly prevalent, it is interesting to consider how “creative” these models can be. From cognitive science, creativity consists of at least two key characteristics: convergent thinking (purposefulness to achieve a given goal) and divergent thinking (adaptability to new environments or constraints) [44].
In this work, we introduce a framework for quantifying LLM creativity that incorporates the two characteristics. This is achieved by (1) Denial Prompting pushes LLMs to come up with more creative solutions to a given problem by incrementally imposing new constraints on the previous solution, compelling LLMs to adopt new strategies, and (2) defining and computing the NeoGauge metric which examines both convergent and divergent thinking in the generated creative responses by LLMs.
We apply the proposed framework on Codeforces problems, a natural data source for collecting human coding solutions. We quantify NeoGauge for various proprietary and open-source models and find that even the most creative model, GPT-4, still falls short of demonstrating human-like creativity. We also experiment with advanced reasoning strategies (MCTS, self-correction, etc.) and observe no significant improvement in creativity. As a by-product of our analysis, we release NeoCoder dataset for reproducing our results on future models. [^1] Our code and dataset are available at https://github.com/JHU-CLSP/NeoCoder
Creative achievement depended not only on the number of alternatives but also the generation of a single high-quality alternative.
  Mark Runco, Critical Creative Process [44]
@@ 1 Introduction
LLMs have shown increasing performance across a wide variety of natural language tasks [13, 26, 33, 36, 39, 40, 42, 58]. However, the degree to which LLMs possess and utilize creativity in solving tasks remains unclear. An automatic method for evaluating LLMs creativity could help developers better understand the emergence of model behaviors, serve as a design objective and pave the path to solving complex real-world problems.
However, despite the importance of machine creativity in establishing artificial general intelligence [9], only a few works have touched upon it [5, 10, 11, 16, 23, 52, 63] because of two major challenges: (1) eliciting diverse and creative generations is difficult [8, 56, 60], and (2) there are no reliable and comprehensive quantitative measurements of LLM creativity. Below, we explicitly illustrate how we address these two challenges.
@@
LLM generations are often repetitive and regurgitating training data [24, 31, 51, 56, 61], making it hard to elicit creative generations. However, we argue that an effective creativity evaluation should be based on the spectrum of maximal creative responses attained from LLMs. Therefore, we introduce Denial Prompting3.1), a prompting method that iteratively “denies” one of the basic tools, techniques, or strategies used in the previous solution (e.g., A: for loops and B: if-else in Fig.1), thereby pushing LLM to think out-of-the-box and elicit creative generations to its fullest extent.
@@
Another challenge in creativity evaluation is to build a reliable and comprehensive quantitative measurement. We propose that such evaluation should be state-aware: adaptive to different contexts, and human-grounded: compares LLM solutions with historical human solutions. According to many cognitive studies, human creativity is viewed as taking place in the interaction with a person, environment, or another model [2, 14, 15, 18, 19, 25]. Similarly, the essence of LLM creativity is also captured from its interaction with the current state (state-aware) and past human knowledge background (human-grounded). This understanding reveals that creativity evaluation should be dynamic, with an individual’s creativity varying under different contexts. For example, in Fig.1, a solution at state $t=0$ probably not be judged at the same creative level as one at state $t=2$, even if they solve the same problem. Because the latter solution is more likely to use novel techniques that humans hardly thought of, such as C: Recursion, to adapt to increasingly challenging constraints.
@@
To address the second challenge, we propose NeoGauge score (§4) which involves (1) verifying the generated solution is correct and follows the given constraints from Denial Prompting (convergent thinking), and (2) checking whether the solution is novel enough that its techniques have never been previously used in human solutions (divergent thinking). This aligned well with the arguments made by Runco [44] that creative achievement depends on both the number of alternative solutions and the generation of high-quality alternatives. By considering both convergent [34, 46, 47, 48] and divergent [22, 25, 53] creative thinking, NeoGauge not only offers a state-aware evaluation at state T but grounds the evaluation in human knowledge through comparing the generated solutions with historical human solutions.
@@
In our experiments, We apply Denial Prompting on Codeforcecs, [^2] https://codeforces.com/problemset a challenging Text-to-Code task where model solutions can be automatically verified and numerous human solutions exist. Specifically, we retrieve 199 latest problems from Codeforces along with 30 human solutions per problem that have successfully passed unit tests. We then run these problems on Denial Prompting to obtain our dataset NeoCoder which consists of original questions with sequences of temporally relevant and increasingly difficult constriants. Examples of NeoCoder are provided in Table 6. We benchmark a broad range of LLMs on NeoCoder and calculate their NeoGauge scores. Additionally, we evaluate four reasoning strategies (MCTS [59], self-correction [45], planning [29], and sampling [12]) on our dataset to study the correlation between augmented machine intelligence and creativity. In summary, our contributions are twofold:
@@
  • We introduce Denial Prompting to elicit creative generations from LLMs and NeoGauge metric to evaluate LLM creativity that follows the two proposed protocols.
@@ 2 Background and Related Works
1 Machine Creativity Evaluation.
@@
While the extensive studies on human creativity from psychological and cognitive science [3, 20, 22, 35, 44, 50, 53], LLM creativity has received little attention. Existing works in studying LLM creativity [16, 52, 64], however, hardly tackle the above two challenges: (1) incentivizing creative generations, and (2) metrics of evaluation are unreliable and incomprehensive.
@@
Tian et al. [52] release a challenging real-world problem dataset to push LLM to think out-of-the-box, but they do not provide an automatic creativity evaluation method built upon their dataset. Additionally, their problems are constructed from a single constraint. In contrast, our Denial Prompting is formulated for multiple iterations of constraint detection and problem refinement, making the generations more creative and providing more states for creativity evaluation. Zhu et al. [64] and Xu et al. [56] design protocols to dynamically generate challenging problems with controllable constraints. However, they mainly focus on accuracy evaluation rather than creativity.
@@
Chakrabarty et al. [10], DeLorenzo et al. [16], and Zhao et al. [63] introduce automatic evaluation pipelines to quantify the four subcomponents of creativity proposed in the Torrance Tests of Creative Thinking [53]: fluency, flexibility, originality, and elaboration. However, the test is originally designed to study human divergent creative thinking (§2) and is unclear whether it applies to machine creativity. Additionally, their evaluation methods hardly meet the above two proposed protocols.
2 Divergent Creative Thinking.
@@
Divergent thinking is a cognitive process that involves exploring a multitude of potential applications for a given set of tools [25]. It typically occurs spontaneously and randomly, leading to numerous possible solutions. Extensive research [3, 22] has been conducted to study divergent creativity, including popular psychometric approaches such as the Unusual Uses Test [22] and Torrance Tests of Creative Thinking [53]. These are designed to let examinees think of as many uses for a (common or unusual) object as possible. The underlying idea of stimulating creative solutions from constrained and unusual settings is also adopted in our Denial Prompting.
@@
Divergent thinking can also be viewed through the lens of $\mathcal{P}$-creativity (Psychological) and $\mathcal{H}$-creativity (Historical) defined by Boden et al. [7]. A valuable idea is $\mathcal{P}$-creative if the person in whose mind it arises could not have come up with it before. Furthermore, a valuable idea is $\mathcal{H}$-creative if it is $\mathcal{P}$-creative and no one else in human history has ever had it before. $\mathcal{P}$-creativity measurement is already pivoted to the nature of Denial Prompting that at each state, we ask LLM to come up with a brand new solution that it has never thought of before by imposing a new constraint. Therefore, we mainly consider $\mathcal{H}$-creativity measurement in our NeoGauge score, where we compare the model-generated solution with a set of collected human solutions to examine if it has ever been proposed in human history (i.e., the ratio of the region out of human solution space in Fig.1). [^3] In our experiment, we collect 30 human-annotated solutions for each problem to approximate the historical human solution space, which we consider sufficient given the high overlapping rate among human solutions. This makes our NeoGauge human-grounded and reflects the novelty from history.
3 Convergent Creative Thinking.
@@
Since the twenty-first century, more researchers have begun to accept the proposition that creative thought involves not merely the generation of many alternative solutions (divergent thinking) but also the identification of new feasible solutions [6, 44]. They frame this problem-solving process as convergent creative thinking and begin to examine how understanding human cognition and convergent thinking might be used to account for creative thought [20, 35, 50]. Several famous cognitive approaches that study the mental representation and process underlying convergent creative thinking [34] involve asking examinees to predict future states from past states using incomplete information [46, 47], or solving the problems as though the counterfactual premises are true [48, 49]. All these tests share certain characteristics, such as always having a single best answer and asking examinees to think in unconventional ways. In our work, besides computing $\mathcal{H}$-creativity for evaluating divergent thinking, we also measure convergent creativity by verifying the feasibility of the generated solution: whether they are correct and following the given constraints. Our NeoGauge metric delivers a more comprehensive evaluation of machine creativity.
@@ 3.1 Denial Prompting: Eliciting Creative Generations from LLMs
Our purpose is to construct a pipeline that iteratively imposes constraints on previous solutions (e.g., disallowing the use of hashmaps) to force more creative solutions. The setup is as follows: we use a highly capable augmentation model $\mathbf{P}_{\text{LM}}$ (e.g. GPT-4) to generate solutions and scrutinize “technique(s)” used in the generated solution, then update problem by imposing the detected technique as a constraint. We repeat this process $t$ times to obtain consecutive $t$ problems with increasingly hard constraints. Fig.2 shows an example with $t=2$. As shown in Algorithm 1, given a reasoning problem $x$ and an initial empty constraint list $\mathcal{C}_{0}=\{\}$, we first let the augmentation model $\mathbf{P}_{\text{LM}}$ to generate an initial solution $y_{1}\sim\mathbf{P}_{\text{LM}}(x)$ via a default problem-solving prompt and conversation history. We then use the same augmentation model $\mathbf{P}_{\text{LM}}$ to detect atomic techniques (e.g., recursion, for loop, hashmaps, etc.), $\mathcal{T}_{1}=\{\tau^{1},\tau^{2},\cdots,\tau^{i}\}$, used in $y_{1}$ to solve $x$ with a technique detection prompt. Then, one technique is randomly sampled $\tau_{1}\sim\mathcal{T}_{1}\setminus\mathcal{C}_{0}$ to ensure it has never been used before as a constraint. Finally, we update the problem $x$ to $x\oplus\tau_{1}$ by imposing it not to use the detected technique $\tau_{1}$ and update constraint list $\mathcal{C}_{0}$ to $\mathcal{C}_{1}=\{\tau_{1}\}$. This is the first iteration of Denial Prompting. We repeat the process to progressively obtain the overall constraint list $C_{t}=\{\tau_{1},\tau_{2},\cdots,\tau_{t}\}$. The prompts for Denial Prompting (including technique detection; used across all experiments) are in Appendix C.
During Denial Prompting, we use a single conversation thread of $\mathbf{P}_{\text{LM}}$ to infer $y_{t}$. Namely, we put $y_{t}\sim\mathbf{P}_{\text{LM}}(x\oplus\tau_{1}\oplus\cdots\oplus\tau_{t-1})$ (line 2 in Algorithm 1) over $t$ in one single thread, such that the model can utilize the trace of previous interactions (including problem statements, constraints, and LLM solutions from each iteration). In practice, we observed incorporating prior interactions in the context improved model generations. Conversely, when detecting solution techniques $\mathcal{T}_{t}\sim\mathbf{P}_{\text{LM}}(y_{t})$ (line 4 in Algorithm 1) we disregard the context from previous conversation rounds to focus the responses solely on the most recent round.
@@ 3.2 NeoCoder Dataset to Support Benchmarking LLM Creativity
Challenge problems.
To construct our creativity benchmark, we compile $n=199$ latest Codeforces problems. We chose problems with a difficulty of 800 (easiest level) since in our preliminary experiments we observed near-random performance on more challenging problems when using well-known open-source models. Furthermore, we selected the recent data to prevent any memorization during pre-training [27].
@@
For each problem, we extract $m=30$ correct human solutions per problem (total of $59$K human solutions). We use human solutions to measure $\mathcal{H}$-creativity of LLM responses.
@@
We also retrieve all test examples provided with each problem (4.5 test examples per problem on average, total of 2.2K test examples). We then perform manual fixes to address any parsing or formatting issues in the collected test examples and ensure that they follow a standardized input-output format. We use these test examples to measure $\mathcal{P}$-creativity or the functional correctness of LLM responses.
@@
Table 2: Proportion of top 5 most common atomic techniques found in the constraint list at each state and human solutions.
Augmentation with Denial Prompting.
We use GPT-4 [37] as the augmentation model $\mathbf{P}_{\text{LM}}$. [^4] We use gpt-4-1106-preview across all experiments, accessed on Dec 2023 through April 2024. We feed retrieved problems to Denial Prompting with maximum $T=5$ to get our dataset NeoCoder. Our dataset consists of pairs $(x,\mathcal{C}_{t}=\{\tau_{1},\tau_{2},\ldots,\tau_{t}\})$, where $x$ represents a problem (programming challenge), and $\mathcal{C}_{t}$ represents the constraints that must be adhered to when solving the problem. This implies that a single programming problem may be associated with various sets of constraints, forming different pairs accordingly.
Statistics for NeoCoder.
Table 2 shows the number of problems $x$ and the number of the associated constraints $|\mathcal{C}_{t}|$. Note that the number of problems decreases for larger number of constraints. This occurs because Denial Prompting may reach a point where it fails to generate any new constraints after a certain number of iterations (i.e., $\mathcal{T}_{t}\setminus\mathcal{C}_{t-1}=\varnothing$ in Algorithm 1). In such a case, we let $\tau_{t}=\varnothing$ and jump to the next iteration $t+1$ without updating the constraint list $\mathcal{C}_{t}=\mathcal{C}_{t-1}$.
We also compare the distribution of the top 5 most common techniques that are discovered from Denial Prompting with the human distribution (Fig.2). It is evident that, without any constraints, models tend to use easy and common techniques (e.g., for-loops) as humans do. However, as more constraints are imposed, the less common but more sophisticated techniques are being employed.

### v2

@@ Abstract
As LLMs become increasingly prevalent, it is interesting to consider how “creative” these models can be. From cognitive science, creativity consists of at least two key characteristics: convergent thinking (purposefulness to achieve a given goal) and divergent thinking (adaptability to explore new environments or constraints) (Runco, 2003). In this work, we introduce a framework for quantifying LLM creativity that incorporates the two design ingredients: (1) We introduce Denial Prompting which pushes LLMs to develop more creative solutions to a given problem by incrementally imposing new constraints on the previous solution, compelling LLMs to adopt new strategies. (2) We define NeoGauge, a metric that quantifies both convergent and divergent thinking in the generated creative responses by LLMs. We test the proposed framework on Codeforces problems, which serve as both a natural dataset for coding tasks and a collection of prior human solutions. We quantify NeoGauge for various proprietary and open-source models and find that even the most creative model, GPT-4, still falls short of demonstrating human-like creativity. We also experiment with advanced reasoning strategies (MCTS, self-correction, etc.) and observe no significant improvement in creativity. As a by-product of our analysis, we release NeoCoder dataset for reproducing our results on future models.111 Our code and data: github.com/JHU-CLSP/NeoCoder
NAACL’25
Benchmarking Language Model Creativity:
A Case Study on Code Generation
Yining Luthanks: Work done at the Johns Hopkins University.ι  Dixuan Wangα  Tianjian Liα  Dongwei Jiangα  Sanjeev Khudanpurα  Meng Jiangι  Daniel Khashabiα ιUniversity of Notre Dame  αJohns Hopkins University
@@ 1 Introduction
Most recent works on LLM creativity evaluation focus on open-ended generation tasks, such as story-writing (Atmakuru et al., 2024; Gómez-Rodríguez and Williams, 2023; Chakrabarty et al., 2024a, b), paper abstract generation (Lu et al., 2024b), and role-play discussion (Lu et al., 2024a). However, the degree to which LLMs possess and utilize creativity for problem-solving remains unclear. An automatic method for evaluating LLMs creativity could help developers better understand the emergence of model behaviors and serve as a design objective in solving complex real-world problems.
However, despite the importance of creativity evaluation in problem-solving, only a few works have touched upon it (DeLorenzo et al., 2024; Tian et al., 2024) because of two major challenges: (1) eliciting diverse and creative solutions is difficult (Bronnec et al., 2024; Xu et al., 2024a; Zhang et al., 2024a), and (2) there are no reliable and comprehensive quantitative measurements of LLM creativity. Below, we explain how we tackle these two challenges for evaluating LLM creativity in problem-solving settings.
@@
LLM generations are often repetitive and regurgitating training data (Holtzman et al., 2019; Kirk et al., 2024; Tevet and Berant, 2021; Xu et al., 2024a; Zhang et al., 2024b), making it hard to elicit creative generations. However, we argue that an effective creativity evaluation method should be based on the spectrum of maximal creative solutions attained from LLMs. Therefore, we introduce Denial Prompting3.1), a prompting method that iteratively “denies” one of the basic tools, techniques, or strategies used in the previous solution (e.g., A: for loops and B: if-else in Figure 1), thereby pushing LLM to think out-of-the-box and elicit creative generations to its fullest extent.
@@
Another challenge in creativity evaluation is to build a reliable and comprehensive quantitative measurement. We propose that such evaluation should be state-aware—adaptive to different contexts, and human-grounded—comparing LLM-generated solutions to historical human solutions. According to many cognitive studies, human creativity is viewed as taking place in the interaction with a person, environment, or another model (Amabile, 1996; Csikszentmihalyi, 1996, 1998; Feldman, 1998; Feldman et al., 1994; Holyoak and Morrison, 2005). Similarly, the essence of LLM creativity should also be captured from its interaction with the current state (state-aware) and past human knowledge background (human-grounded). This understanding reveals that creativity evaluation should be dynamic, with an individual’s creativity varying under different contexts. For example, in Figure 1, a solution at state $t=0$ probably will not be judged at the same creative level as one at state $t=2$, even if they solve the same problem. Because the latter solution is more likely to use novel techniques that humans hardly thought of, such as C: Recursion, to adapt to increasingly challenging constraints.
@@
To address the second challenge, we propose NeoGauge score (§4) which involves (1) verifying the correctness of the LLM-generated solution and whether it adheres to the specified constraints from Denial Prompting (convergent thinking), and (2) assessing solution novelty by contrasting it with techniques previously used in human solutions (divergent thinking). This aligns well with the arguments made by Runco (2003) that creative achievement depends on both the number of alternative solutions and the generation of high-quality alternatives. By considering both convergent (Lubart, 2001; Sternberg, 1981, 1982; Sternberg and Gastel, 1989a) and divergent (Guilford, 1950; Holyoak and Morrison, 2005; Torrance, 1966) creative thinking, NeoGauge not only offers a state-aware evaluation but grounds the evaluation in collective human knowledge through comparing the generated solutions with historical human solutions.
@@
In our experiments, we apply Denial Prompting on Codeforces, [^2] https://codeforces.com/problemset a challenging Text-to-Code task where model solutions can be automatically verified and allows comparison to substantial historical human solutions. [^3] We provide detailed justifications for task choice in §A. Specifically, we retrieve 199 latest problems from Codeforces along with 30 human solutions per problem that have successfully passed unit tests. We then run these problems on Denial Prompting to obtain our dataset NeoCoder which consists of original questions with sequences of temporally relevant and increasingly difficult constraints. Examples of NeoCoder are provided in Table 4. We benchmark a broad range of LLMs on NeoCoder and calculate their NeoGauge scores. Additionally, we evaluate four reasoning strategies, MCTS (Zhang et al., 2023), self-correction (Shinn et al., 2023), planning (Jiang et al., 2023b), and sampling (Chen et al., 2021), on our dataset to study the correlation between augmented machine intelligence and creativity. In summary, our contributions are twofold:
@@
  • We introduce Denial Prompting to elicit creative generations from LLMs and NeoGauge metric to evaluate LLM creativity in problem-solving that follows the two proposed protocols.
@@ 2 Background and Related Works
Machine Creativity Evaluation.
@@
While the extensive studies on human creativity from psychological and cognitive science (Amabile, 1982; Finke et al., 1996; Guilford, 1950; Mumford et al., 1991; Runco, 2003; Sternberg and Lubart, 1991; Torrance, 1966), LLM creativity has received little attention. Existing works studying LLM creativity in problem-solving settings (DeLorenzo et al., 2024; Tian et al., 2024; Zhu et al., 2024), however, tend to overlook two challenges: (1) eliciting creative LLM solutions, and (2) ensuring evaluation metrics are grounded and comprehensive.
@@
Tian et al. (2024) have released a challenging real-world problem dataset to push LLM to think out-of-the-box, but they do not provide an automatic creativity evaluation method built upon their dataset. Additionally, their problems are constructed from a single constraint. In contrast, our Denial Prompting is formulated for multiple iterations of constraint detection and problem refinement, making the generations more creative and providing more states for creativity evaluation. Another concurrent work (Atmakuru et al., 2024) also employs multiple constraints to facilitate creative generation; however, their evaluation primarily targets linguistic creativity (Lu et al., 2024b) and it is tested on open-ended story writing task. Zhu et al. (2024) and Xu et al. (2024a) design protocols to dynamically generate challenging problems with controllable constraints. However, their evaluation mainly focuses on accuracy rather than creativity.
@@
Chakrabarty et al. (2024a), DeLorenzo et al. (2024), and Zhao et al. (2024) introduce automatic evaluation pipelines to quantify the four subcomponents of creativity proposed in the Torrance Tests of Creative Thinking (Torrance, 1966): fluency, flexibility, originality, and elaboration. However, the test is originally designed to study human divergent creative thinking (§2) and is unclear whether it applies to machine creativity.
Divergent Creative Thinking.
@@
Divergent thinking is a cognitive process that involves exploring a multitude of potential applications for a given set of tools (Holyoak and Morrison, 2005). It typically occurs spontaneously and randomly, leading to numerous possible solutions. Extensive research (Amabile, 1982; Guilford, 1950) has been conducted to study divergent creativity, including popular psychometric approaches such as the Unusual Uses Test (Guilford, 1950). These are designed to let examinees think of as many uses for a (common or unusual) object as possible. The underlying idea of stimulating creative solutions from constrained and unusual settings is also adopted in our Denial Prompting.
@@
Divergent thinking can also be viewed through the lens of $\mathcal{P}$-creativity (Psychological) and $\mathcal{H}$-creativity (Historical) defined by Boden et al. (1994). A valuable idea is $\mathcal{P}$-creative if the person in whose mind it arises could not have come up with it before. Furthermore, a valuable idea is $\mathcal{H}$-creative if it is $\mathcal{P}$-creative, and no one else in human history has ever had it before. $\mathcal{P}$-creativity measurement is embedded in the structure of Denial Prompting, where at each state, the LLM is prompted to come up with a brand new solution that it has never thought of before by imposing a new constraint. Therefore, we mainly consider $\mathcal{H}$-creativity measurement in our NeoGauge score, where we compare the model-generated solution with a set of collected human solutions to examine if it has ever been proposed in human history (i.e., the ratio of the region out of human solution space in Figure 1). This makes our NeoGauge human-grounded and reflects the novelty from history.
Convergent Creative Thinking.
@@
Since the twenty-first century, more researchers have begun to accept the proposition that creative thought involves not merely the generation of many alternative solutions (divergent thinking) but also the identification of new feasible solutions (Baer, 1994; Runco, 2003). They frame this problem-solving process as convergent creative thinking and begin to examine how understanding human cognition and convergent thinking might be used to account for creative thought (Finke et al., 1996; Mumford et al., 1991; Sternberg and Lubart, 1991). Several famous cognitive approaches that study the mental representation and process underlying convergent creative thinking (Lubart, 2001) involve asking examinees to predict future states from past states using incomplete information (Sternberg, 1981, 1982), or solving the problems as though the counterfactual premises are true (Sternberg and Gastel, 1989a, b). All these tests share certain characteristics, such as always having a single best answer and asking examinees to think in unconventional ways. In our work, besides computing $\mathcal{H}$-creativity for evaluating divergent thinking, our work also measures convergent creativity by verifying the feasibility of the generated solution: whether they are correct and following the given constraints. Our NeoGauge metric delivers a more comprehensive evaluation of machine creativity.
@@ 3.1 Denial Prompting: Eliciting Creative Generations from LLMs
Our purpose is to construct a pipeline that iteratively imposes constraints on previous solutions (e.g., disallowing the use of hashmaps) to force more creative solutions. The setup is as follows: given an input problem, we use a highly capable augmentation model $\mathbf{P}_{\text{LM}}$ (e.g. GPT-4) to generate solutions and scrutinize “technique(s)” used in the generated solution, then update the problem by imposing the detected technique as a constraint. We repeat this process $t$ times to obtain consecutive $t$ problems with increasingly hard constraints (Figure 8 shows an example with $t=2$).
Specifically, as shown in Algorithm 1, given a reasoning problem $x$ and an initial empty constraint list $\mathcal{C}_{0}=\{\}$, we first let the augmentation model $\mathbf{P}_{\text{LM}}$ to generate an initial solution $y_{1}\sim\mathbf{P}_{\text{LM}}(x)$ via a default problem-solving prompt and conversation history. We then use the same augmentation model $\mathbf{P}_{\text{LM}}$ to detect atomic techniques (e.g., recursion, for loop, hashmaps, etc.), $\mathcal{T}_{1}=\{\tau^{1},\tau^{2},\cdots,\tau^{i}\}$, used in $y_{1}$ to solve $x$ with a technique detection prompt. Then, one technique is randomly sampled $\tau_{1}\sim\mathcal{T}_{1}\setminus\mathcal{C}_{0}$ to ensure it has never been used before as a constraint. Finally, we update the problem $x$ to $x\oplus\tau_{1}$ which explicitly prohibits the use of the technique $\tau_{1}$ and update constraint list $\mathcal{C}_{0}$ to $\mathcal{C}_{1}=\{\tau_{1}\}$. [^4] We use $\oplus$ to indicate text concatenation. This is the first iteration of Denial Prompting. We repeat the process to progressively obtain the overall constraint list $C_{t}=\{\tau_{1},\tau_{2},\cdots,\tau_{t}\}$. The prompts for Denial Prompting (including technique detection; used across all experiments) are in Appendix D.
@@
During Denial Prompting, we use a single conversation thread of $\mathbf{P}_{\text{LM}}$ to infer $y_{t}$ such that the model can utilize the trace of previous interactions (including problem statements, constraints, and LLM solutions from each iteration). In practice, we observe adding prior interactions in the context improves model generations. Conversely, when detecting solution techniques $\mathcal{T}_{t}\sim\mathbf{P}_{\text{LM}}(y_{t})$ (line 4 in Algorithm 1), we disregard the context from previous conversation rounds to focus the responses solely on the most recent round.
@@ 3.2 NeoCoder Dataset to Support Benchmarking LLM Creativity
Challenging problems.
To construct our creativity benchmark, we compile $n=199$ latest Codeforces problems. We chose problems with a difficulty of 800 (easiest level) since, in our preliminary experiments, we observed near-random performance on more challenging problems when using well-known open-source models. Furthermore, we selected the recent data to prevent any memorization during pre-training Huang et al. (2023).
@@
For each problem, we extract $m=30$ correct human solutions per problem (total of $5.9$K human solutions). [^5] We consider 30 human-annotated solutions to construct a historical solution space for each problem to be sufficient given the high overlap rate among them. We use human solutions to measure $\mathcal{H}$-creativity of LLM responses.
@@
We also retrieve all test examples provided with each problem (4.5 test examples per problem on average, a total of 2.2K test examples). We then perform manual fixes to address any parsing or formatting issues in the collected test examples and ensure that follow a standardized input-output format. We use these test examples to measure $\mathcal{P}$-creativity or the functional correctness of LLM responses.
Augmentation with Denial Prompting.
We use GPT-4 (OpenAI, 2024) as the augmentation model $\mathbf{P}_{\text{LM}}$ because we find that GPT-4 can achieve 94% technique detection recall compared to the human programmer in our pilot experiments. [^6] We use gpt-4-1106-preview across all experiments, accessed from Dec 2023 through April 2024. We feed the retrieved problems to Denial Prompting3.1) with maximum iterations $T=5$ to obtain our dataset NeoCoder. Our dataset consists of pairs $(x,\mathcal{C}_{t}=\{\tau_{1},\tau_{2},\ldots,\tau_{t}\})$, where $x$ represents a problem (programming challenge), and $\mathcal{C}_{t}$ represents the constraints that must be adhered to when solving the problem $x$. This implies that a single programming problem may be associated with various sets of constraints, forming different pairs accordingly.
Statistics for NeoCoder.
Table 1 shows the number of problems $x$ and the number of the associated constraints $|\mathcal{C}_{t}|$. Note that the number of problems decreases for a larger number of constraints. This is due to Denial Prompting potentially reaching a point where it can no longer generate new constraints after a certain number of iterations (i.e., $\mathcal{T}_{t}\setminus\mathcal{C}_{t-1}=\varnothing$ in Alg. 1). In such a case, we let $\tau_{t}=\varnothing$ and jump to the next iteration $t+1$ without updating the constraint list $\mathcal{C}_{t}=\mathcal{C}_{t-1}$.
@@
We also compare the distribution of the top 5 most common techniques from Denial Prompting in comparison to that of human solutions (Figure 2). It is evident that, without any constraints, models tend to use common techniques (e.g., for-loops) similar to human solutions. However, as more constraints are imposed, the less common but more sophisticated techniques are employed.
Figure 2: Proportion of the top 5 most common atomic techniques used by GPT-4 per state, compared to those in human solutions. In absense of any constraints (the first column), the model default to common and accessible techniques, like humans (the last column). This echoes our claim in §1 that eliciting creative solutions is crucial for creativity evaluation.